# Towards a Multi-Embodied Grasping Agent

### v1

@@ V-A Geometric Scene Encoder
The scene geometry is represented by a point cloud $\{\mathbf{r}_{i}\}_{i=1}^{N}$, typically obtained from a dense RGB-D scan. Our method relies exclusively on geometric information from the point cloud. This design choice is motivated by the goal of robust sim-to-real transfer, as it circumvents the significant challenges of photorealistic rendering and domain randomization required to align simulated RGB data with real-world conditions. While color could be included as a straightforward scalar feature, our geometry-only approach ensures greater robustness out of the box. The raw point cloud is first pre-processed via voxel down-sampling (1 mm grid size) and then mapped to a fixed cardinality of 15,000 points using the Farthest-Point Sampling (FPS) algorithm. This fixed-size representation is necessary to accommodate JAX’s [10] requirement for statically-sized arrays.
Our encoder constructs a hierarchical representation of the scene through a feature pyramid composed of $L$ stages. At the initial stage, we compute an equivariant feature embedding $\mathbf{f}_{i}^{(0)}$ for each point $\mathbf{r}_{i}$ by aggregating information from its local neighborhood $\mathcal{N}(i)$ as a weighted sum
$\mathbf{f}_{i}^{(0)}=\sum_{j\in\mathcal{N}(i)}w_{ij}\ \mathrm{dir\_enc}\bigl(\hat{\mathbf{r}}_{ji},r_{ji}\bigr).$
Here, $\mathrm{dir\_enc}$ encodes the relative direction $\hat{\mathbf{r}}_{ji}$ and distance $r_{ji}$ from a neighbor point $\mathbf{r}_{j}$ to the center point $\mathbf{r}_{i}$. The weights $w_{ij}$ are distance-based attention scores, learned via the expression $w_{ij}=\mathrm{softmax}_{j}(\mathrm{MLP}(r_{ji}))$.
Each subsequent pooling stage $k\in\{1,\dots,L\}$ downsamples the feature point cloud $\{(\mathbf{r}_{i}^{(k-1)},\mathbf{f}_{i}^{(k-1)})\}$ using FPS to produce a sparser set of points $\{\mathbf{r}_{i}^{(k)}\}$. The features for these new points are computed by aggregating features from the previous, denser layer using the sparse directional SO(2) tensor product operation in the EquiformerV2 modules. This process creates a multi-scale geometric representation $\{(\mathbf{r}_{i}^{(k)},\mathbf{f}_{i}^{(k)})\}_{k=0..L}$. To expand the receptive field at each level, we apply multiple equivariant message-passing layers, where the neighborhood graphs are constructed using the K-Nearest Neighbors (K-NN) algorithm to ensure a fixed number of neighbors per point. This hierarchical structure enables efficient, point-wise feature querying by concatenating features across all stages for a given spatial location. The design is inspired by the Sonata model [64], which addresses the geometric shortcut problem present in U-Net-style architectures. By eliminating the decoder path, our approach significantly reduces memory utilization compared to similar equivariant models [9, 5], while preserving a rich, multi-scale geometric representation.
Figure 3: Equivariant Gripper Embeddings. An initial gripper configuration (a) is represented by a learned feature embedding $z$. After a physical joint rotation $\Delta R$, the gripper is in a new configuration (b). Our method ensures the feature is correspondingly transformed via the Wigner-D matrix, $z^{\prime}=\mathbf{D}^{\ell}(\Delta R)z$, keeping the representation consistent with the physical state.
@@ V-B Geometric Gripper Encoding
The gripper encoder’s objective is to produce a set of configuration-aware, equivariant query features that represent the gripper’s state. A naive encoding of the scalar joint values $\mathbf{q}_{g}$ would force the model to learn the gripper’s kinematic structure implicitly, hindering generalization. To avoid this, we adopt a more explicit, geometry-aware approach.
For each joint $i$, we introduce a unique, learnable, equivariant feature embedding, $\mathbf{z}_{i}$. Each embedding is composed of multiple irreducible representations (irreps), e.g., $\ell_{0}\oplus\ell_{1}\oplus\ell_{2}$. At a minimum, we use scalar ($\ell=0$) and vector ($\ell=1$) components. The gripper’s forward kinematics, $\mathrm{kin}_{g}(\mathbf{q}_{g},i)=T_{i}=(R_{i},\mathbf{p}_{i})$, are used to place and orient these features in the gripper’s base frame. Specifically, for each joint, the rotation $R_{i}$ is applied to its corresponding embedding $\mathbf{z}_{i}$ via the Wigner-D matrices
$\mathbf{z}^{\prime}_{i}=\bigoplus_{\ell}D^{\ell}(R_{i})\mathbf{z}_{i}^{\ell}.$
@@
This fundamental transformation, which ensures the feature representation covaries with the physical state, is visualized in Figure 3. This operation aligns the feature embedding with the joint’s local orientation. To recover the scalar features of a revolute joint, we leverage the kinematic chain. We compute a channel-wise dot product between the rotated child embedding $\mathbf{z}^{\prime}_{i}$ and its parent’s rotated embedding, producing channel-wise scalars that are a function of the relative rotation between features. These scalars are then used in a lightweight attention mechanism to modulate the channels of $\mathbf{z}^{\prime}_{i}$ with the dir_mod operator, thereby conditioning it on the whole gripper configuration. This scalar mechanism only requires the relative transform between parent and child, making it robust to the choice of coordinate systems for each gripper. For joints at the root of a kinematic chain, a static auxiliary embedding is used as an anchor for this computation, which is discarded afterwards. To incorporate information from prismatic joints, we apply an equivariant message-passing layer over the kinematic graph. The messages utilize the dir_mod operation on the relative positions $(\mathbf{p}_{i}-\mathbf{p}_{j})$, which encodes the translational component of the joints. Finally, to represent the gripper’s 6-DoF global pose, we introduce two additional learnable embeddings anchored at the gripper’s origin. In the decoding phase, the $\ell=1$ components of these two embeddings are used to predict the rotational and translational velocity fields (the twist). Crucially, this representation circumvents the challenge of canonical coordinate frame selection, a problem present in prior works such as DiffusionEDFs [5] and [1]. Consequently, the local frames of different grippers are not required to be manually pre-aligned; their kinematic origins are sufficient. The learnable nature of the embeddings allows the model to implicitly establish a consistent internal frame and modulate the various frequency components of the output rotational flow. This process yields a total of $D_{g}+2$ query features, each with an associated position, that explicitly encode the gripper’s kinematic structure and configuration. This flexibility makes the equivariant design not just a feature for improving data efficiency, but a fundamental requirement for our multi-embodiment approach.
@@ V-C Gripper-Scene Relation
The Multiscale Tensor Field module is responsible for relating the gripper query features to the encoded scene geometry. This step provides the necessary context for predicting the gripper’s pose relative to the scene and for adapting the finger placement. First, the set of $D_{g}+2$ gripper query features, along with their associated positions, are transformed from the gripper’s local frame into the global scene frame based on the current SE(3) pose estimate in the flow process.
We then construct a bipartite graph to establish neighborhoods. For each query, we identify its K-nearest neighbors among the scene points at the highest-resolution stage of the feature pyramid. Leveraging the many-to-one mapping established during the encoding phase, we then retrieve the corresponding features for these neighbors from all levels of the hierarchy. These hierarchical scene features are first projected to a lower-dimensional latent space and then conditioned on the flow-time $t$ using an equivariant FiLM layer [1].
For each edge connecting a query to a scene, we apply the edge-direction-dependent $\mathrm{SO}(2)$ tensor product [62] on the scene feature. This operation is conditioned on a set of scalars derived from the flow-time, the relative distance, and the dot products of the query and scene features. The resulting edge features from a query’s neighborhood are then aggregated into a single, scene-context feature using an attention mechanism. This aggregated scene feature is then fused with the original gripper query feature using a final FCTP. Note, the expensive FCTP operations are only applied per query thus limiting the computational and memory impact. This entire query-relation block can be stacked multiple times to build a more detailed representation. Finally, the scene-aware query features are transformed back into the gripper’s local frame for decoding.
@@ V-D Flow Decoding
After the gripper queries have been enriched with scene context, we apply several additional equivariant message-passing layers over the gripper’s kinematic graph. The final flow vector is then decoded by extracting specific components from the feature embeddings: the $\ell=1$ components of the two pose tokens yield the rotational velocity $\boldsymbol{\omega}$ and the translational velocity $\mathbf{v}$, while the $\ell=0$ components of the $D_{g}$ joint tokens yield the joint-space velocities $\dot{\mathbf{q}}$. The model is trained by minimizing a weighted mean squared error between the predicted flow $(\boldsymbol{\omega},\mathbf{v},\dot{\mathbf{q}})$ and its ground-truth counterpart, following the flow-matching framework of Lim et al. [37].
VI Experimental Results
TABLE I: Parallel Jaw Grasp success rate comparison
Method Grasp Success Rate (%)
OrbitGrasp 96.7
Ours 94.8
Grasp success rate on the OrbitGrasp [6] benchmark. Models are trained on 2 500 scenes. While only competitive, our method only uses successful grasps during training, operates on full scenes and does not require scene normals.
We compare our method with the state-of-the-art grasping method OrbitGrasp [6] and in a multi-embodiment grasping settings  [^2] The results presented for our method are preliminary, as they were obtained from models that had not yet fully converged (both single and multi-embodiment models), and are therefore subject to change.. It is worth noting that, compared with the previously mentioned methods, ours does not prioritize grasp success rate but rather reproducing the distribution of successful grasps. This distinction provides flexibility in adapting our method to the desired downstream task with minimal modifications.
@@ VI-A Single Embodiment Parallel Jaw Grasping
First, we evaluate the feasibility of our approach on the benchmark presented in OrbitGrasp [6] for SE(3) grasp pose prediction of the Panda gripper in cluttered object scenes. For a direct comparison with OrbitGrasp [6], we reproduced its original dataset, which comprises 2,500 cluttered ‘pile’ scenes containing approximately 2.5 million successful and 3.5 million unsuccessful grasps. Adhering to the original evaluation protocol, we assess the grasp success rate in a simulated de-cluttering procedure across 100 scenes. The results are presented in Table I  [^3] Preliminary results of our model. The OrbitGrasp result is from the original publication [6]. Final evaluation results are subject to change. For the OrbitGrasp model, we select the variant with irreps up to $\ell=2$, as this aligns with our model’s architecture and has near negligible performance differences to significantly more computationally expensive versions [6].
The results show that while OrbitGrasp nearly saturates the success rate metric, our method remains competitive. Notably, our model achieves this performance despite being trained exclusively on successful grasps—utilizing less than half of the available grasp data—and without requiring scene normal or color information as input. Furthermore, there are two key methodological differences. First, our approach processes the full, high-resolution scene point cloud (15,000 points) at once, in contrast to OrbitGrasp which operates on multiple low-resolution scene patches of fewer than 1,000 points each. Second, the OrbitGrasp architecture is constrained to predicting per-point grasp scores along the object’s surface normal. This inherent design choice limits its applicability to more general datasets, such as ours, where grasp poses can be freely positioned within the scene.
@@
Gripper Total Grasps
Franka 6,870,351
VX300s 6,238,898
DEX-$E$ 1,639,770
Allegro 2,116,060
Shadow 2,990,167
Available grasps per gripper over all scenes. For each gripper we generate 5 000 scenes and re-evaluate all stable gravity free grasps per object in scene.
@@ VI-B Dataset Generation
Our evaluation requires a dataset composed of scenes containing grasps with explicit joint configurations for each gripper type. The data generation pipeline from MultiDiffusion [1] is unsuitable for our purposes, as it only provides grasps in fixed open and closed configurations and employs the Antipodal [15] method, which is not designed for generating diverse, high-DoF grasps. For our dataset, we follow a two-stage process. First, we generate stable grasps for each gripper-object pair in a gravity-free environment. Second, these candidate grasps are transferred to cluttered scenes, where they are filtered for collisions and their stability is re-evaluated via a lifting motion, similar to the protocol in [1].
@@
For the Panda and VX300s parallel-jaw grippers, we employ the Antipodal [15] grasp generation method to collect 5,000 grasps per object. For the dexterous grippers (DEX-EE, Allegro, and Shadow Hand), we leverage a differentiable kinematics model implemented in JAX [10], identical to the one used in our main method. We define a set of potential contact points on each gripper’s fingertips. For each object, we first sample 10 cm grasping regions by applying FPS to the object’s surface points. Within each region, we randomly select target contact points on the object and corresponding target points on the gripper’s fingers; object contact points are slightly offset from the surface to mitigate initial collisions. We then solve for the corresponding joint configurations via gradient descent by minimizing an objective function. The primary term is a positional consistency loss,
$L_{\text{pos}}=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\mathbf{p}_{\text{finger}}^{i}-\mathbf{p}_{\text{target}}^{i}\right\rVert^{2},$
where $\mathbf{p}_{\text{finger}}^{i}$ denotes the position of the $i$-th fingertip and $\mathbf{p}_{\text{target}}^{i}$ is its assigned target point. The secondary term encourages stable contacts by penalizing the angular deviation between surface normals, defined as
@@
$L_{\text{norm}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\cos(\theta_{i})\right),$
@@
where $\theta_{i}$ is the angle between the normals at the $i$-th contact pair. Although this optimization in some cases yields a stable grasp in only one to two percent of attempts, its key advantage is its generality; a single, parallelizable method can synthesize a large quantity of grasps in a single batch for any dexterous gripper. Grasp stability is validated in the MuJoCo [65] physics simulator. A candidate grasp is first checked for collisions. If collision-free, the gripper is closed, and a series of impulse forces are applied to the object. A grasp is deemed stable if the gripper maintains contact throughout all checks. This impulse-based procedure is significantly faster than the shaking-motion approach used in [1], while empirically showing more stable grasps. Finally, the validated stable grasps are placed back into the cluttered scenes, following the final data processing steps of [1]. Table II summarizes the total number of grasps generated for each gripper. Due to their geometric complexity and higher degrees of freedom, the dexterous grippers consequently have a lower final grasp density per scene in our dataset. In total, our generated dataset encompasses about 22 500 scenes with over 200 objects sourced from the Google Scanned Objects (GSO) [66] and YCB [67] dataset. Each scene contains one to seven objects and includes up to 1 500 grasps.
@@
TABLE III: Multi-embodiment grasp success rates across various grippers.
Method Panda Franka (2 DoF) ViperX 300s (2 DoF) DEX-EE (12 DoF) Allegro (16 DoF) Shadow Hand (22 DoF)
Single-Embodiment 93.9 83.7 72.1 73.6 72.4
Multi-Embodiment 97.0 88.3 74.5 66.1 79.6
@@
Grasp success percentage per gripper evaluated over 10 scenes testing 100 sampled grasps. Training data set for each gripper encompasses 5 000 scenes. Single-Embodiment models are trained on only gripper specific data, while Multi-Embodiment model is trained on all scenes combined.

### v2

@@ V-A Geometric Scene Encoder
The scene geometry is represented by a point cloud $\{\mathbf{r}_{i}\}_{i=1}^{N}$, typically obtained from dense RGB-D. While color could be included as a scalar feature, our method relies exclusively on geometric information, avoiding the challenges of photorealistic rendering and domain randomization required to align simulated RGB data with real-world conditions. The raw point cloud is voxel down-sampled (1 mm grid) and mapped to a fixed cardinality of 15,000 points using Farthest-Point Sampling (FPS). This fixed size accommodates JAX’s [9] static array requirement.
Our encoder constructs a hierarchical representation through a feature pyramid of $L$ stages. At the initial stage, we compute an equivariant feature encoding $\mathbf{f}_{i}^{(0)}$ for each point $\mathbf{r}_{i}$ by aggregating information from its local neighborhood $\mathcal{N}(i)$ as a weighted sum $\mathbf{f}_{i}^{(0)}=\sum_{j\in\mathcal{N}(i)}w_{ij}\ \mathrm{dir\_mod}\bigl(\hat{\mathbf{r}}_{ji},r_{ji}\bigr).$ Here, $\mathrm{dir\_mod}$ encodes the relative direction $\hat{\mathbf{r}}_{ji}$ and distance $r_{ji}$ from a neighbor $\mathbf{r}_{j}$ to the center $\mathbf{r}_{i}$. The weights $w_{ij}$ are distance-based attention scores, learned via $w_{ij}=\mathrm{softmax}_{j}(\mathrm{MLP}(r_{ji}))$.
Each subsequent pooling stage $k\in\{1,\dots,L\}$ downsamples the cloud $\{(\mathbf{r}_{i}^{(k-1)},\mathbf{f}_{i}^{(k-1)})\}$ using FPS to produce a sparser set $\{\mathbf{r}_{i}^{(k)}\}$. Features for these points are computed by aggregating from the previous denser layer using the sparse directional SO(2) tensor product in EquiformerV2 modules. This creates a multi-scale representation $\{(\mathbf{r}_{i}^{(k)},\mathbf{f}_{i}^{(k)})\}_{k=0..L}$. To expand the receptive field, we apply multiple equivariant message-passing layers with K-Nearest Neighbors (K-NN) graphs. This hierarchical structure enables efficient, point-wise feature querying by concatenating features across all stages. Inspired by Sonata [66], which addresses the geometric shortcut problem in U-Nets, we eliminate the decoder path. This significantly reduces memory utilization compared to similar equivariant models [8, 4], while preserving a rich, multi-scale representation.
@@ V-B Geometric Gripper Encoding
The gripper encoder produces configuration-aware, equivariant query features representing the gripper’s state. A naive encoding of scalar joint values $\mathbf{q}_{g}$ would force implicit learning of kinematics, hindering generalization. We adopt an explicit, geometry-aware approach.
@@
For each joint $i$, we introduce a unique, learnable, equivariant feature embedding $\mathbf{z}_{i}$, composed of multiple irreducible representations (irreps), e.g., $\ell_{0}\oplus\ell_{1}\oplus\ell_{2}$. At minimum, we use scalar ($\ell=0$) and vector ($\ell=1$) components. The forward kinematics $\mathrm{kin}_{g}(\mathbf{q}_{g},i)=T_{i}=(R_{i},\mathbf{p}_{i})$ place and orient these features in the gripper’s base frame. Specifically, rotation $R_{i}$ is applied to its embedding $\mathbf{z}_{i}$ via Wigner-D matrices: $\mathbf{z}^{\prime}_{i}=\bigoplus_{\ell}D^{\ell}(R_{i})\mathbf{z}_{i}^{\ell}$. This transformation ensures the representation covaries with the physical state (visualized in Figure 1), aligning the embedding with the joint’s local orientation. To recover scalar features of a revolute joint, we leverage the kinematic chain. We compute a channel-wise dot product between the rotated child embedding $\mathbf{z}^{\prime}_{i}$ and its parent’s rotated embedding, producing scalars as a function of the relative rotation. These scalars modulate the channels of $\mathbf{z}^{\prime}_{i}$ via a lightweight attention mechanism and the dir_mod operator, conditioning it on the gripper configuration. This mechanism only requires the relative transform between parent and child, making it robust to coordinate system choices. For root joints, a static auxiliary embedding serves as an anchor and is discarded afterwards. To incorporate prismatic joints, we apply an equivariant message-passing layer over the kinematic graph. Messages utilize dir_mod on relative positions $(\mathbf{p}_{i}-\mathbf{p}_{j})$, encoding the translational component.
Finally, to represent the gripper’s 6-DoF global pose, we introduce two additional learnable embeddings anchored at the gripper origin. In the decoding phase, the $\ell=1$ components of these predict the rotational and translational velocity fields (the twist). Crucially, this circumvents canonical coordinate frame selection challenges present in works such as DiffusionEDFs [4] and [10]. Local frames of different grippers need not be manually pre-aligned; kinematic origins suffice. The learnable embeddings allow the model to implicitly establish a consistent internal frame. This yields $D_{g}+2$ query features, each with an associated position, explicitly encoding the gripper structure. This flexibility makes the equivariant design a fundamental requirement for our multi-embodiment approach.
@@ V-C Gripper-Scene Relation
The Multiscale Tensor Field module relates gripper query features to encoded scene geometry, providing context for predicting gripper pose and adapting finger placement. First, the $D_{g}+2$ gripper query features and their positions are transformed from the local frame to the global scene frame based on the current SE(3) pose estimate.
We construct a bipartite graph to establish neighborhoods. For each query, we identify K-nearest neighbors among scene points at the highest-resolution pyramid stage. Leveraging the many-to-one mapping from encoding, we retrieve corresponding features from all hierarchy levels. These hierarchical features are projected to a lower-dimensional latent space and conditioned on flow-time $t$ using an equivariant FiLM layer [10].
For each edge connecting a query to a scene point, we apply the edge-direction-dependent $\mathrm{SO}(2)$ tensor product [64] on the scene feature. This is conditioned on scalars derived from flow-time, relative distance, and dot products of query and scene features. Resulting edge features are aggregated into a scene-context feature using attention, then fused with the original gripper query using a final FCTP. Note that expensive FCTP operations are only applied per query, limiting computational impact. This block can be stacked to build detailed representations. Finally, scene-aware query features transform back into the gripper’s local frame for decoding.
@@ V-D Flow Decoding
After enriching gripper queries with scene context, we apply additional equivariant message-passing layers over the kinematic graph. The final flow is decoded by extracting specific components: $\ell=1$ components of the two pose tokens yield rotational velocity $\boldsymbol{\omega}$ and translational velocity $\mathbf{v}$, while $\ell=0$ components of the $D_{g}$ joint tokens yield joint-space velocities $\dot{\mathbf{q}}$. The model is trained by minimizing a weighted mean squared error between predicted flow $(\boldsymbol{\omega},\mathbf{v},\dot{\mathbf{q}})$ and ground-truth samples, following common flow-matching frameworks [37, 39, 38, 40].
TABLE I: Parallel-jaw grasp performance: Success and declutter rate
Method Success Rate (%) Declutter Rate (%)
OrbitGrasp [5] 98.4 98.2
Ours 93.7 96.2
Grasp success and declutter rates and on the OrbitGrasp [5] benchmark over 100 scenes. Models are trained on 2 500 scenes. While competitive, our method uses less than half the grasps during training and applies no post-processing.
VI Experiments
We compare our method with the state-of-the-art grasping method OrbitGrasp [5] and in multi-embodiment grasping settings. It is worth noting that, compared with the previously mentioned methods, ours does not prioritize grasp success rate but rather reproducing the distribution of successful grasps. This distinction provides flexibility in adapting our method to the desired downstream task with minimal modifications.
@@ VI-A Single Embodiment Parallel Jaw Grasping
First, we evaluate the feasibility of our approach on the benchmark presented in OrbitGrasp [5] for SE(3) grasp pose prediction of the Panda gripper in cluttered object scenes. For a direct comparison with OrbitGrasp [5], we reproduced its original dataset, which comprises 2,500 cluttered ‘pile’ scenes containing approximately 2.5 million successful and 3.5 million unsuccessful grasps. Adhering to the original evaluation protocol, we assess the grasp success rate and declutter rate in a simulated decluttering procedure across 100 scenes. The results are presented in Table I. For the OrbitGrasp model, we select the variant with irreps up to $\ell=2$, as this aligns with our model’s architecture and has near-negligible performance differences to significantly more computationally expensive versions with higher $\ell$-types [5].
The results show that while OrbitGrasp [5] nearly saturates the success rate metric, our method remains competitive. Notably, our model achieves this performance despite being trained exclusively on successful grasps—utilizing less than half of the available grasp data—and without requiring scene normal or color information as input. Furthermore, there are two key methodological differences. First, our approach processes the full, high-resolution scene point cloud (15,000 points) at once, in contrast to OrbitGrasp which operates on multiple low-resolution scene patches of fewer than 1,000 points each. Second, the OrbitGrasp architecture is constrained to predicting per-point grasp scores along the object’s surface normal. This inherent design choice limits its applicability to more general datasets, such as ours, where grasp poses can be freely positioned within the scene.
@@
Franka ViperX DEX-$E$ Allegro Shadow
6.87M 6.24M 1.58M 2.12M 2.99M
Total available grasps (in millions) per gripper over all scenes. For each gripper we generate 5 000 scenes and re-evaluate all stable gravity-free grasps per object in the scene.
@@ VI-B Dataset Generation
TABLE III: Grasp success rate comparison for static grippers
Method Franka DEX-$E$ Shadow
Diffusion [10] 97.1 93.5 85.8
Ours 97.7 84.8 91.0
Grasp success rate comparison of methods trained on static grippers, following Diffusion [10], for a direct comparison. We only predict flow and diffusion for the pose and fix the joint configurations.
@@
Our evaluation requires a dataset composed of scenes containing grasps with explicit joint configurations for each gripper type. The data generation pipeline from Diffusion [10] is unsuitable for our purposes, as it only provides grasps in fixed open and closed configurations and employs the Antipodal [15] method, which is not designed for generating diverse, high-DoF grasps. For our dataset, we follow a two-stage process. First, we generate stable grasps for each gripper-object pair in a gravity-free environment. Second, these candidate grasps are transferred to cluttered scenes, where they are filtered for collisions and their stability is re-evaluated via a lifting motion, similar to the original protocol [10].
@@
For the Panda and ViperX parallel-jaw grippers, we employ the Antipodal [15] grasp generation method to collect 5,000 grasps per object. For the dexterous grippers (DEX-EE, Allegro, and Shadow Hand), we leverage a differentiable kinematics model implemented in JAX [9], identical to the one used in our main method. We define a set of potential contact points on each gripper’s fingertips. For each object, we first sample 10 cm grasping regions by applying FPS to the object’s surface points. Within each region, we randomly select target contact points on the object and corresponding target points on the gripper’s fingers; object contact points are slightly offset from the surface to mitigate initial collisions. We then solve for the corresponding joint configurations via gradient descent by minimizing an objective function.The loss terms consist of a positional consistency loss and an angular deviation between surface normals,
@@
$L_{\text{pos}}=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\mathbf{p}_{\text{f}}^{i}-\mathbf{p}_{\text{t}}^{i}\right\rVert^{2},\quad L_{\text{norm}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\cos(\theta_{i})\right),$
@@
where $\mathbf{p}_{\text{f}}^{i}$ denotes the position of the $i$-th fingertip, $\mathbf{p}_{\text{t}}^{i}$ is its assigned target point, and $\theta_{i}$ is the angle between the normals at the $i$-th contact pair. Although this optimization, in some cases, yields a stable grasp in only one to two percent of attempts, its key advantage is its generality; a single, parallelizable method can synthesize a large quantity of grasps in a single batch for any dexterous gripper. Grasp stability is validated in the MuJoCo [67] physics simulator. A candidate grasp is first checked for collisions. If collision-free, the gripper is closed, and a series of impulse forces are applied to the object. A grasp is deemed stable if the gripper maintains contact throughout all checks. This impulse-based procedure is significantly faster than the shaking-motion approach used in [10], while empirically showing more stable grasps. Finally, the validated stable grasps are placed back into the cluttered scenes, following the final data processing steps of [10]. Table II summarizes the total number of grasps generated for each gripper. Due to their geometric complexity and higher degrees of freedom, the dexterous grippers consequently have a lower final grasp density per scene in our dataset. In total, our generated dataset encompasses 25 000 scenes with over 200 objects sourced from the Google Scanned Objects (GSO) [68] and YCB [69] datasets. Each scene contains one to seven objects and includes up to 1 500 grasps.
@@
TABLE IV: Multi-embodiment grasp success rates, coverage, and joint variance.
Franka (2 DoF) ViperX (2 DoF) DEX-EE (12 DoF) Allegro (16 DoF) Shadow (22 DoF)
Method Succ. NJD. Succ. NJD. Succ. NJD. Succ. NJD. Succ. NJD.
Single-Embodiment 98.1 0.249 98.3 0.148 80.4 0.110 82.9 0.281 79.3 0.218
Multi-Embodiment 97.9 0.247 95.9 0.172 74.4 0.115 86.2 0.295 80.7 0.252
@@
Grasp success percentage (Succ.) and mean normalized joint deviation (NJD.) per gripper evaluated over 10 scenes testing 100 sampled grasps. Training dataset for each gripper encompasses 5 000 scenes.